OpenStudy (anonymous):

consider line "a" that passes through (14,2) and (28,6) write an equation for line "a".

OpenStudy (saifoo.khan):

Step 1: FInd the slope. Step 2: use, \[y-y_1 = m(x-x_1) \]To find the equation.

OpenStudy (anonymous):

\[y-y _{0}=m(x-x _{0})\] m= 6-2/28-14 just substitute\[ y _{0} ,x_{0}\]

OpenStudy (anonymous):

y-2=2/7(x-14) rearanging you can put it in y=mx +b form

OpenStudy (anonymous):

(y-y1)/(x-x1)=(y2-y1)/(x2-x1) (y-2)/(x-14)=(6-2)/(28-14) (y-2)/(x-14)=4/14 y-2=2(x-14)/7 Solving the above will give you the equation of line a .

OpenStudy (anonymous):

(y2-y1)/(x2-x1) -> this is the slope m of the line